### v1

# ReflecSched: Solving Dynamic Flexible Job-Shop Scheduling via LLM-Powered Hierarchical Reflection

Abstract
Dynamic Flexible Job-Shop Scheduling (DFJSP) is an NP-hard problem challenged by real-time event adaptation and complex machine routing. While traditional dispatching rules are efficient but rigid, deep learning approaches are opaque and require intricate feature engineering. Large Language Models (LLMs) promise adaptive reasoning without this engineering overhead, yet we find their direct application is suboptimal. Baseline LLMs suffer from three key pitfalls: the long-context paradox, where crucial data is underutilized; an underutilization of expert heuristics; and myopic decision-making. To address this, we propose ReflecSched, a framework that empowers the LLM beyond a direct scheduler by equipping it with a strategic analysis capability. ReflecSched tasks the LLM to analyze heuristic-driven simulations across multiple planning horizons and distill them into a concise, natural-language summary termed “Strategic Experience”. This summary is then integrated into the prompt of a final decision-making module, guiding it to produce non-myopic actions. Experiments show that ReflecSched not only statistically significantly outperforms direct LLM baselines, securing a 71.35% Win Rate and a 2.755% Relative Percentage Deviation reduction, but also surpasses the performance of all individual heuristics evaluated, all while demonstrably mitigating the three identified pitfalls. Additionally, ReflecSched performs on par with the best heuristic tailored to each instance across all problem cases.
1 Introduction
Dynamic Flexible Job-Shop Scheduling (DFJSP) is a long-standing challenge in operations research and a key component of modern smart manufacturing (Li et al. 2025). As an NP-hard problem, it involves the complex task of continuously allocating operations to machines in a real-time environment characterized by stochastic events such as new job arrivals or machine breakdowns. The ability to efficiently schedule these operations critically impacts the agility, resilience, and profitability of production systems. For decades, the field has been dominated by methods ranging from classical heuristics to meta-heuristics (Cao et al. 2023). While foundational, these approaches often rely on handcrafted rules that can struggle to generalize across diverse and unforeseen dynamic scenarios (Ferreira, Figueira, and Amorim 2022).
In recent years, deep learning (DL), particularly reinforcement learning (RL), has emerged as a promising paradigm for this problem (Xu et al. 2025). By training agents on simulated data, DL-based methods can learn sophisticated, state-aware scheduling policies. However, this capability comes at a significant cost. A key limitation lies in the intricate process of state and action space engineering: translating the rich state of a dynamic workshop into a fixed-size numerical vector that a neural network can process is a complex task requiring dual expertise in scheduling and machine learning (Zhang et al. 2020). Furthermore, the resulting models are often “black boxes” with opaque decision-making logic that is difficult for engineers to interpret or trust in high-stakes environments (Ngwu, Liu, and Wu 2025).
The advent of Large Language Models (LLMs) offers a different approach, potentially circumventing these challenges by leveraging natural language as a more accessible interface. An LLM-based scheduler could, in principle, reason directly over a textual description of the factory state, reducing the need for complex numerical encoding (Abgaryan, Harutyunyan, and Cazenave 2024). This suggests the possibility of developing and adapting scheduling strategies with greater flexibility.
However, our work finds that a direct application of LLMs to DFJSP leads to consistently suboptimal performance. We argue that the standard autoregressive generation process is not inherently well-suited for the strategic, lookahead-dependent reasoning that scheduling demands. Through a motivational analysis, we identify and empirically validate three key pitfalls of this direct approach:
  • The Long-Context Paradox: Foundational static information, such as machine processing times and job structures, is often underutilized by the model as the prompt length increases.
  • Underutilization of Heuristics: LLMs exhibit difficulty in reliably applying expert-provided procedural knowledge, such as Priority Dispatching Rules (PDRs), often allowing their generalized pre-trained behaviors to override explicit guidance.
  • Myopic Greed: The lack of a structured planning mechanism creates a strong tendency towards locally optimal but globally inefficient decisions, which can lead to downstream bottlenecks.
To address these shortcomings, we introduce ReflecSched, a framework that fundamentally restructures the LLM’s role in the scheduling process. Rather than acting as a purely reactive decision-maker, the LLM also serves as a strategic analyst within a planning loop that decouples long-horizon reasoning from immediate execution. At its core, ReflecSched’s Hierarchical Reflection module performs multi-level, heuristic-driven simulations of future trajectories. It then tasks the LLM to reflect upon these simulations and distill them into a concise, actionable “Strategic Experience.” This experience, generated only at critical, event-driven moments, subsequently guides a lean, fast Experience-Guided Decision-Making module to produce a high-quality, non-myopic action.
Our contributions are threefold:
  1. 1.
    We present a systematic investigation into the application of modern LLMs to the full scope of the DFJSP, incorporating a rich set of dynamic events.
  2. 2.
    We provide an empirical analysis that identifies and validates three key failure modes of baseline LLM-based schedulers, offering insights for the broader field of LLM-based planning.
  3. 3.
    We propose ReflecSched, a framework that addresses these pitfalls through a combination of hierarchical simulation and LLM-driven experience synthesis.
2 Related Work
Heuristic and Metaheuristic Approaches.
Heuristic-based methods are a foundational approach for the Dynamic Job-Shop Scheduling Problem (DJSP) due to their computational tractability. Research has advanced from simple dispatching rules to complex metaheuristics designed to handle diverse operational constraints. For instance, some works have applied GRASP algorithms to manage a wide array of dynamic events (Baykasoğlu, Madenoğlu, and Hamzadayı 2020), while others have used hybrid Particle Swarm Optimization for joint production and transportation scheduling (Ren et al. 2022). Further research has focused on specific uncertainties, such as variable processing times, using techniques like Artificial Bee Colony algorithms (Shahgholi Zadeh, Katebi, and Doniavi 2019). A key limitation of these metaheuristics, however, is their reliance on a pre-defined search logic that is inherently less adaptive and struggles to generalize to new, unseen problem scenarios.
Deep Reinforcement Learning Approaches.
To address the generalization limitations of heuristics, Deep Reinforcement Learning (DRL) has emerged as an alternative that learns adaptive policies, typically by formulating the DJSP as a Markov Decision Process. Architectural innovations have been central to this area. For example, Pointer Networks have been utilized for end-to-end policy learning that avoids reliance on predefined rules (Yang et al. 2024). To enhance scalability, a multi-agent perspective has been adopted through decentralized frameworks (Liu, Piplani, and Toro 2023). Other approaches decompose the decision-making process using hierarchical structures (Luo, Zhang, and Fan 2021). Nevertheless, DRL presents its own significant challenges: performance is highly sensitive to the complex engineering of state and reward functions; the resulting policies are often opaque “black boxes” hindering trust; and they demand accurate simulation environments for training, which are often unavailable for complex real-world systems.
Figure 1: An example of two valid schedules for the same problem instance. The Gantt charts illustrate how a locally optimal choice at an early stage can lead to a suboptimal final makespan (top, 7.51), whereas a more globally-aware decision sequence yields a better outcome (bottom, 6.51).
Large Language Models Paradigms.
Seeking to address the challenges of both handcrafted logic and the black-box nature of DRL, LLMs have recently introduced a new paradigm for scheduling by leveraging their vast pre-existing knowledge and reasoning capabilities. One direction explores LLMs as end-to-end solvers by fine-tuning them on large-scale datasets of problems paired with their solutions (Abgaryan, Cazenave, and Harutyunyan 2025; Abgaryan, Harutyunyan, and Cazenave 2024). Another approach employs LLMs as automated problem modelers that translate natural language into executable code for traditional solvers (Amarasinghe et al. 2023; AhmadiTeshnizi et al. 2024). A third direction positions the LLM as a hyper-heuristic generator that iteratively reflects upon and improves heuristic rule-code (Huang et al. 2024). Despite their potential, early works reveal critical challenges: the risk of generating infeasible solutions (“hallucinations”), high computational costs, and a primary focus on static problems, leaving their application to complex, dynamic scheduling environments largely unexplored.
3 Preliminaries
Dynamic Flexible Job-Shop Scheduling.
DFJSP is the problem of assigning operations from multiple jobs to a set of machines, subject to real-time stochastic events. It is characterized by two features: flexibility, which allows each operation to be processed on a subset of candidate machines, and dynamism, which refers to the occurrence of real-time stochastic events such as new job arrivals or machine breakdowns. The objective is to generate a valid schedule that minimizes the makespan, defined as the maximum completion time among all jobs (Cao and Yuan 2024). Figure 1 illustrates an example of such a schedule, visualized as a Gantt chart. A valid schedule must satisfy both precedence constraints between operations of the same job and the resource constraints of the machines. A formal mathematical formulation is presented in Appendix A.
Baseline LLM-based Scheduler.
To evaluate our proposed method, we define a direct baseline, LLM-Direct, which employs the LLM to perform the function of classical dispatching rules. In this approach, a state-aware prompt is constructed at each decision point. The shop-floor state is encoded into the prompt, including both static specifications, such as job structures and machine data, and dynamic information, such as the set of currently available actions and current machine statuses. The action selected by the LLM is then executed in a shop-floor simulator, advancing the system state to the next decision point. Further implementation details for LLM-Direct are presented in Appendix B.
4 Motivational Analysis: The Pitfalls of Direct LLM-based Scheduling
4.1 The Long-Context Paradox
While replacing state engineering with a single LLM prompt is appealing, it creates an information utility paradox: encoding the high-dimensional shop-floor state into a dense, verbose prompt leads to underutilization of the provided data (Jiang et al. 2024). We observe this is not merely a context-length issue but a fundamental challenge in reasoning over such textual complexity (Liu et al. 2023), as exemplified by our case study with smaller models (Appendix C).
To validate this hypothesis, we conducted a targeted experiment by omitting the static information block from the LLM-Direct baseline’s prompt. As shown in Figure 2, this omission had a negligible impact on performance, even though the static block constitutes over 70% of the prompt. The median makespan ratio between the conditions (with vs. without static data) is nearly 1.0, providing strong evidence that the model largely ignores this information and fails to ground its decisions in the provided specifications.
(a) Prompt Composition
(b) Performance Impact
Figure 2: An empirical investigation of the Long-Context Paradox. (a) A breakdown of the prompt composition for the baseline model. (b) Box plots showing the ratio of makespans resulting from running the model with and without the static data portion of the prompt.
4.2 Underutilization of Heuristics
A key challenge in LLM-based scheduling is the effective application of PDRs. While LLMs are generally proficient at following declarative instructions, their capacity to faithfully execute complex, state-dependent procedural rules remains unreliable (Mu et al. 2024).
To isolate and measure this capability, we developed PDR-Bench, a diagnostic benchmark where each instance is programmatically constructed to have a single, known optimal heuristic, and employed it in a controlled experiment to evaluate our baseline LLM (see Appendix D for detailed experimental details). The “optimal heuristic” for any given instance in PDR-Bench is known by design, as each problem instance is specifically constructed to ensure that one particular heuristic yields a superior outcome over all others. The results (Figure 3) reveal the LLM’s significant difficulty in effectively utilizing these heuristics (Uzunoglu, Safa, and Şahin 2024). This finding strongly suggests the model’s inability to translate the strategic guidance of the provided rules into its decision-making. Furthermore, when explicitly instructed to use the single optimal heuristic, the model still failed to match the performance of that heuristic applied in isolation, reverting instead to its generalized, pre-trained behaviors.
Figure 3: Analysis of heuristic utilization on the PDR-Bench dataset, with Relative Percent Deviation (RPD) measured against the known optimal heuristic for each instance (0% baseline). The baseline LLM struggles to apply the optimal heuristic even when explicitly prompted, whereas ReflecSched substantially improves performance.
Figure 4: The architecture of the ReflecSched framework, designed to decouple strategic planning from immediate execution. The main workflow is initiated at a decision point. The Hierarchical Reflection Module (blue) performs multi-level, heuristic-driven simulations to explore future state trajectories. It then leverages an LLM to analyze these outcomes and distill a concise Strategic Experience ($\mathcal{E}$). This experience is passed to the Experience-Guided Decision-Making Module (green), which constructs an experience-augmented prompt to guide the LLM in selecting a final, strategically-informed action. The right-hand panels provide concrete examples of the prompts and responses for both the reflection and decision-making stages, illustrating the flow of information from simulation data to strategic guidance, and finally to a specific action.
4.3 The Pitfall of Myopic Greed
The autoregressive, token-by-token generation process inherent in LLMs presents a structural mismatch with the strategic, non-local search required for complex scheduling problems. This mismatch predisposes them to myopic decision-making, where locally optimal choices lead to globally suboptimal outcomes.
Figure 5: Comparison of the Average Greedy Decision Ratio (GDR) between the LLM-Direct baseline and the ReflecSched framework. Results are presented for different models and are segmented by the Normal and Small problem scales.
To quantify this behavior, we evaluated baseline LLMs using the Greedy Decision Ratio (GDR) on GEN-Bench, a benchmark specifically designed to control for structural biases that could otherwise mask or mimic such tendencies. The quantitative results (Figure 5) reveal that the GDR is persistently high across all tested models, indicating a significant greedy bias (Baeumel, van Genabith, and Ostermann 2025). A qualitative case study (Figure 1) further illustrates the consequences, showing how a single myopic choice can trigger a cascade of downstream inefficiencies. This combination of quantitative and qualitative data demonstrates that myopic behavior, stemming from the models’ autoregressive nature, is a critical pitfall for direct generation approaches in scheduling.
5 The ReflecSched Framework
5.1 Framework Overview
To address the identified limitations, we introduce ReflecSched, a novel framework that structurally decouples long-horizon, strategic reasoning from low-latency, online decision-making via two interconnected, sequential modules (Figure 4) (Yao et al. 2023b). The first, the Hierarchical Reflection Module, serves as the framework’s strategic planning component. It leverages multi-level, heuristic-driven simulations to explore future state trajectories (Yao et al. 2023a). It then employs the LLM to analyze these simulated outcomes, distilling them into a concise, actionable Strategic Experience $\mathcal{E}$. This experience then informs the second component, the Experience-Guided Decision-Making Module, which handles real-time execution. Guided by the strategic foresight in $\mathcal{E}$, this module makes a strategically-informed decision based on the immediate environment state.
This two-stage architecture is designed to address each of the identified challenges: First, by explicitly simulating future consequences, the Hierarchical Reflection Module provides the necessary foresight to counteract myopic greed, shifting the model’s focus from locally optimal choices to globally efficient strategies. Second, the framework directly addresses the underutilization of heuristics by employing PDRs to guide its exploratory rollouts (Zhang et al. 2024). This ensures that expert procedural knowledge is integrated into the reasoning process. Finally, the framework mitigates the long-context paradox. By having the Reflection Module distill its findings into a concise Strategic Experience $\mathcal{E}$, the final decision prompt supplied to the LLM contains only this high-level guidance and the immediate state (Jiang et al. 2024). This precludes informational overload and concentrates the model’s reasoning on the most salient data.
5.2 Hierarchical Reflection
Central to ReflecSched is the Hierarchical Reflection module, a mechanism designed to perform strategic evaluation, which is essential for the DFJSP. It implements a recursive “Simulate-Reflect-Refine” loop (Gui et al. 2025). This iterative process first explores the future through simulation, then reflects on the outcomes to identify key strategic principles, and finally refines these principles into a concise, actionable policy directive (Shinn et al. 2023).
The process is a top-down hierarchical simulation from $l_{max}$ to $l=0$, operating as a deterministic projection that assumes no new stochastic events during the lookahead. Central to this simulation is a randomized base policy, $\pi_{base}$, which operates by sampling a PDR from a predefined pool at each decision point to select an action. The simulation methodology differs by level to enable multi-timescale exploration. At higher levels ($l>0$), the module performs sparse, long-range exploration by simulating trajectories where $\pi_{base}$ is applied at every step. In contrast, the base level ($l=0$) performs a systematic evaluation of immediate actions. For each available action $a\in A(S_{t})$, a short-horizon rollout is initiated by first executing the action $a$, and then applying $\pi_{base}$ for all subsequent decisions.
The outcome of any such rollout is a trajectory $\tau^{(l)}$ and its associated cost, defined as the hypothetical partial makespan:
$\hat{J}^{(l)}(S_{t})=\text{Makespan}(S_{t+\text{steps}(l)})$ (1)
Following the simulation at a given level $l$, the reflection phase distills actionable insights from the raw trajectory data. This approach is designed to clearly delineate the attributes of successful versus unsuccessful strategies by focusing on extremal outcomes. Specifically, the module identifies the trajectories that resulted in the minimum and maximum cost estimates from the set of simulations performed at that level:
$\tau^{(l)}_{best}$ $=\arg\min_{\tau^{(l)}\text{ from }S_{t}}\hat{J}^{(l)}(S_{t})$ (2)
$\tau^{(l)}_{worst}$ $=\arg\max_{\tau^{(l)}\text{ from }S_{t}}\hat{J}^{(l)}(S_{t})$ (3)
Finally, in the refinement phase, these two contrasting trajectories are provided to the LLM, which acts as a higher-level policy synthesizer (Light et al. 2025). Its task is not merely to summarize but to identify the key distinguishing features between $\tau^{(l)}_{best}$ and $\tau^{(l)}_{worst}$ and distill these observations into a concise, textual strategic guideline, termed the Strategic Experience $\mathcal{E}$. This synthesis, represented by the function $F_{LLM}$, effectively transforms high-dimensional, numerical simulation data into low-dimensional, human-interpretable strategic guidance:
$\mathcal{E}=F_{LLM}\left(\tau^{(l)}_{best},\tau^{(l)}_{worst}\right)$ (4)
To ensure operational efficiency, this computationally intensive reflection process is invoked selectively. It is triggered only upon the occurrence of a dynamic event (e.g., a new job arrival or machine breakdown), as such an event may render the existing Strategic Experience $\mathcal{E}$ suboptimal or obsolete.
5.3 Experience-Guided Decision-Making
The strategic experience $\mathcal{E}$ distilled by the reflection module provides the long-horizon context that LLM-Direct schedulers lack. This module translates this high-level guidance into a specific, executable action by constructing a concise, experience-augmented prompt. This prompt, containing only the immediate dynamic state and the guidance from $\mathcal{E}$, is then used to query the LLM for a final decision.
Theoretically, this process connects to the principles of Approximate Policy Iteration (API) (Gordon 1995). The validity of this connection, and the resulting performance guarantee, hinges on two key working assumptions regarding the reflection process: (1) the rollout cost $\hat{J}$ serves as an approximation of the true cost-to-go function for the base policy $\pi_{\mathrm{base}}$; and (2) the LLM operates as a faithful synthesizer, meaning its generated experience $\mathcal{E}$ correctly reflects the superiority of the best simulated trajectory over the worst (Faithful Reflection).
The experience $\mathcal{E}_{t}$ generated under these conditions is then used to define the final policy $\pi_{\mathcal{E}_{t}}$, which makes a greedy decision with respect to the guidance (Schmied et al. 2025):
$a^{\star}=\pi_{\mathcal{E}_{t}}(S_{t})$ (5)
With the policy thus defined, our framework inherits the classical policy-improvement guarantee of rollout algorithms (Bertsekas 2012):
Theorem 1 (Monotonic Improvement Guarantee).
Let $\pi_{\mathcal{E}_{t}}$ be the policy defined above. If the assumptions of cost function approximation and Faithful Reflection hold, then for all states $S_{t}$, its expected performance is no worse than that of the base heuristic policy, $\pi_{\mathrm{base}}$:
$V^{\pi_{\mathcal{E}_{t}}}(S_{t})\;\leq\;V^{\pi_{\mathrm{base}}}(S_{t})$ (6)
Figure 6: Comparative performance of ReflecSched and the LLM-Direct baseline on the GEN-Bench dataset. Bars (left y-axis) represent the average Relative Percent Deviation (RPD), where lower is better. The green line (right y-axis) shows the Win Rate of ReflecSched against the baseline. For reference, the performance of the single best-performing PDR (i.e., the heuristic with the lowest average RPD across all instances) and the average of all PDRs are shown as red and purple dashed lines, respectively. P-values from a Wilcoxon signed-rank test are annotated above each RPD bar group.
Theorem 1 establishes a conditional theoretical performance guarantee, grounding ReflecSched in established reinforcement learning principles and ensuring that, under plausible assumptions, its performance will not degrade below that of the underlying heuristics. While a formal proof of the underlying assumptions is intractable for black-box LLMs, our empirical study Section 6 shows a consistent and significant reduction in makespan. This suggests that these assumptions hold sufficiently well in practice for performance improvements to be realized. This combination of a conditional theoretical guarantee and strong empirical performance suggests that ReflecSched is an effective and well-grounded framework for dynamic scheduling.
6 Experiments
We conduct a series of experiments to evaluate the ReflecSched framework. In Section 6.1, we detail the experimental setup, which includes the benchmark datasets, evaluation metrics, hyperparameter tuning process, and the specific protocols used for different analyses. In Section 6.2, we present a comparative performance analysis to assess ReflecSched’s effectiveness against both the LLM-Direct baseline and traditional heuristics, and to empirically verify its mitigation of the identified pitfalls. In Section 6.3, we conduct an ablation study to isolate and quantify the contribution of the framework’s core hierarchical reflection mechanism.
6.1 Experimental Setup
Benchmark Datasets.
Our evaluation is performed on two benchmark suites we developed for this study: GEN-Bench, designed for general-purpose evaluation, and PDR-Bench, a diagnostic benchmark specifically constructed to assess heuristic utilization. Both suites contain instances at two scales: Normal (20 instances for GEN-Bench, 112 for PDR-Bench) and Small (20 for GEN-Bench, 12 for PDR-Bench). The detailed generation and curation process for these datasets is described in Appendix E.
Metrics and Models.
Our primary metrics are: (1) Relative Percent Deviation (RPD) (Huang, Gao, and Li 2025) calculated against the best makespan found for each instance across all methods and all runs evaluated in this study and (2) Win Rate, the percentage of instances where ReflecSched achieves a strictly better makespan than LLM-Direct. To assess the statistical significance of the performance difference, we use the non-parametric Wilcoxon signed-rank test (Demšar 2006). We evaluate a range of LLMs, including GPT-4o (Hurst et al. 2024), DeepSeek-V3 (Liu et al. 2024), and the Qwen3 series (8B, 14B, 32B) (Yang et al. 2025). All models are prompted in a zero-shot manner, utilizing Chain-of-Thought reasoning (Wei et al. 2022). The specific prompt structures, which encode shop-floor state and strategic guidance into natural language, are detailed in Appendix F.
Hyperparameter Tuning.
The core hyperparameters for ReflecSched are its reflection depth, Level ($L$), and its exploration breadth, the Number of Rollouts ($R$). To determine a suitable configuration, we conducted a systematic analysis on a dedicated validation set of 40 instances using the Qwen3-8B model, with average RPD as the performance metric. Based on this analysis, we selected the configuration of $L=6$ and $R=24$, which yielded the best average RPD on the validation set, for all subsequent experiments. Detailed setup and results of the hyperparameter analysis are provided in Appendix G.
Evaluation Protocol.
Our evaluation protocol is adapted to the specific goal of each experiment. For the initial diagnostic analysis (Sections 4.1-4.3), our aim is to capture the baseline model’s intrinsic behavioral tendencies. To achieve this, we employ a multi-sample voting scheme: actions are decided by the majority vote from 5 independent generations at a higher temperature ($T=0.8$) (Wang et al. 2022). In contrast, for the final comparative experiments designed to assess practical performance, all LLM inferences within the ReflecSched framework and for the LLM-Direct baseline use a low temperature ($T=0.2$) for efficient and stable decision-making. To mitigate outcome variability due to model stochasticity and provide a more stable performance estimate, we report the average results across 3 independent runs.
6.2 Comparative Performance Analysis
Our experimental goal is to primarily investigate the effectiveness of our hierarchical reflection architecture. Therefore, our main comparison is against the LLM-Direct baseline, with strong heuristics included as a classical performance reference. A direct comparison to DRL methods is omitted due to fundamental differences in methodology, as our work is motivated by circumventing their complex state engineering requirements.
Instance Set RPD (%) $p$-value
Normal 0.0620 0.8107
Small 0.9331 0.1240
Table 1: Statistical comparison between ReflecSched and the best-performing heuristic on each instance across all problem cases.
ReflecSched demonstrates a statistically significant performance improvement over the LLM-Direct baseline across all tested configurations (Wilcoxon signed-rank test, $p<0.05$), securing a 71.35% average Win Rate and an average RPD reduction of 2.755% (see Figure 6 for details). Furthermore, the framework demonstrates the capability to outperform all evaluated PDR, with several of the tested LLM backbones enabling performance superior to any individual heuristic. For a further comparison against a baseline constructed by selecting the best-performing heuristic for each instance, the resulting average RPDs are minimal, and a two-sided Wilcoxon signed-rank test confirms no statistically significant difference between their performances (Table 1, $p>0.05$).
The framework’s superior performance can be explained by its success in mitigating the specific pitfalls identified in our analysis. First, it addresses the underutilization of heuristics, with results on the diagnostic PDR-Bench (Figure 3) showing that ReflecSched consistently and substantially reduces the RPD compared to the baseline across all evaluated models. Second, it directly counteracts myopic greed; the data in Figure 5 shows a consistent reduction in the GDR across all models, providing strong evidence that the generated strategic experience steers the LLM towards non-myopic decisions. Finally, it circumvents the long-context paradox. By having the reflection module distill complex state information into a concise experience $\mathcal{E}$, the final decision prompt becomes substantially shorter and more focused. This is empirically validated by our token consumption analysis (Appendix H), which shows that on the larger, more complex Normal-scale instances, ReflecSched requires fewer tokens than the baseline. This efficiency gain is less pronounced on simpler Small-scale problems, where the benefits of such strategic distillation offer naturally diminished returns.
Figure 7: Ablation study on the hierarchical reflection mechanism. The chart compares the average RPD of the full ReflecSched framework with that of a single-level version. The performance of the single-level model is shown across a range of different values for the Number of Rollouts.
6.3 Ablation Study
To isolate the contribution of our hierarchical reflection architecture, we conducted an ablation study comparing the full ReflecSched framework ($L=6,R=24$) against an ablated, single-level version ($L=0$). To determine if increased computational effort at a single level could compensate for the lack of hierarchical depth, this ablated model was evaluated across a range of search breadths $R\in\{1,3,6,12,24\}$.
The results presented in Figure 7 unequivocally demonstrate the superiority of the full ReflecSched framework. Across all tested search breadths, the full framework consistently outperforms the single-level model—whose performance, even at its maximum computational budget ($R=24$), remains substantially worse.
The results yield two key insights. The performance gap confirms that the strategic foresight gained from multi-level reflection is a primary driver of the framework’s effectiveness. More revealingly, simply increasing the search breadth for the single-level model fails to improve its performance, which plateaus and can even degrade with more rollouts. This strongly suggests that a brute-force, ”flat” search is ineffective without hierarchical guidance, and confirms that the hierarchical reflection mechanism provides a qualitative benefit that simply increasing computation cannot replicate.
7 Conclusion
This paper first provided an empirical diagnosis of applying LLMs to DFJSP, identifying three critical pitfalls: the long-context paradox, heuristic underutilization, and myopic decision-making. To address these, our proposed ReflecSched framework recasts the LLM as a strategic analyst, distilling insights from multi-level simulations into a “Strategic Experience” to guide execution. Experiments confirm that ReflecSched significantly outperforms direct LLM baselines, surpasses all traditional heuristics overall, and maintains performance on par with the instance-specific best heuristic across all problem instances. We posit that the core principle of decoupling strategic reflection from execution offers a promising template for applying LLMs to a broader class of sequential decision-making problems.